# Evaluating Search-Based Software Microbenchmark Prioritization

### v2

@@ 7.2 Software Performance Testing
Performance testing with microbenchmarks is a relatively new area of research. To this end, earlier studies investigate unit-level performance testing on OSS Java projects [59, 76], which conclude that it is still a niche technique, especially compared to unit testing. Explanations could be that more effort is required to write microbenchmarks, the extra cost of their execution, and the higher complexity required to assess the results with statistical analyses. Samoaa and Leitner [75] investigate parameterization of benchmarks in detail. Grambow et al. [31, 32] utilize information from application benchmarks to guide microbenchmark execution and assess the ability of microbenchmarks to detect application performance changes, with limited success. He et al. [37] and Laaber et al. [54] take an orthogonal approach to optimizing performance testing: they dynamically stop benchmarks when their results are of a desired statistical quality instead of optimizing the execution order of the benchmark suite. Traini et al. [79] further assess whether dynamically stopping is beneficially for measuring in steady state.
There exist works on performance regression testing, i.e., greedy heuristics [17], genetic algorithms [3, 4], and machine learning [11] have been used to select the optimal benchmarks for a given software version in terms of their ability to find performance changes. These works focus on RTS whereas this paper targets SMBP. In addition, some works exist for performance regression testing of concurrent classes (e.g., [70, 83]). They focus on concurrent software and generate tests for these, whereas this paper optimizes the execution of existing benchmark suites. Finally, Huang et al. [40] proposed a prediction method to assess the need for performance testing of a new software version, which targets the version level compared to ours on the benchmark level.
We consider two works on performance regression testing closely related to ours [67, 56]. For a given software version, Mostafa et al. [67] prioritize test cases with complex performance impact analysis that requires measuring individual components of not only the SUT but also the Java Development Kit (JDK). Their technique focuses on collection-intensive software and is partially only evaluated on unit tests used as performance tests, whereas our technique is generally applicable to any kind of software and is evaluated only on benchmarks. Laaber et al. [56] studied coverage-based, greedy SMBP techniques and found that they are only marginally more effective than a random ordering. We build on their work and use SAs and novel Greedy techniques instead.
@@ 8 Conclusions and Future Work
This paper defines SBSMBP and Greedy SMBP techniques relying on three objectives: coverage, coverage overlap among benchmarks, and performance change size obtained from historical data. With an extensive experimental evaluation on $10$ Java projects with JMH suites, we study the effectiveness and efficiency of a total of $18$ new SMBP techniques and compare these to two coverage-based, greedy SMBP baselines. The results reveal that the best SBSMBP technique is competitive with the best greedy baselines but does not improve on its effectiveness. Surprisingly, the Greedy technique relying only on the historical performance change size can be more effective than the best coverage-based greedy baselines and SBSMBP techniques while being significantly more efficient.

### v4

@@ 7.2 Software Performance Testing
Performance testing with microbenchmarks is a relatively new area of research. To this end, earlier studies investigate unit-level performance testing of OSS Java projects [59, 76], which conclude that it is still a niche technique, especially compared to unit testing. Explanations could be that more effort is required to write microbenchmarks, the extra cost of their execution, and the higher complexity required to assess the results with statistical analyses. Samoaa and Leitner [75] investigate parameterization of benchmarks in detail. Grambow et al. [31, 32] utilize information from application benchmarks to guide microbenchmark execution and assess the ability of microbenchmarks to detect application performance changes, with limited success. He et al. [37] and Laaber et al. [54] take an orthogonal approach to optimizing performance testing: they dynamically stop benchmarks when their results are of a desired statistical quality instead of optimizing the execution order of the benchmark suite. Traini et al. [79] further assess whether dynamically stopping is beneficially for measuring in steady state.
There exist works on performance regression testing, i.e., greedy heuristics [17], genetic algorithms [3, 4], and machine learning [11], which have been used to select the optimal benchmarks for a given software version in terms of their ability to find performance changes. These works focus on RTS whereas this paper targets SMBP. In addition, some works exist for performance regression testing of concurrent classes (e.g., [70, 83]). They focus on and generate tests for concurrent software, whereas this paper optimizes the execution of existing benchmark suites. Finally, Huang et al. [40] proposed a prediction method to assess the need for performance testing of a new software version, which targets the version level compared to ours on the benchmark level.
We consider two works on performance regression testing closely related to ours [67, 56]. For a given version, Mostafa et al. [67] prioritize test cases with complex performance impact analyses that require measuring individual components of not only the SUT but also the Java Development Kit (JDK). Their technique focuses on collection-intensive software and is partially only evaluated on unit tests used as performance tests, whereas our technique is generally applicable to any kind of software and is evaluated only on benchmarks. Laaber et al. [56] studied coverage-based, greedy SMBP techniques and found that they are only marginally more effective than a random ordering. We build on their work and use SAs and novel Greedy techniques instead.
@@ 8 Conclusions and Future Work
This paper defines SBSMBP and Greedy SMBP techniques relying on three objectives: coverage, coverage overlap among benchmarks, and performance change size obtained from historical data. With an extensive experimental evaluation on $10$ Java projects with JMH suites, we study the effectiveness and efficiency of a total of $18$ new SMBP techniques and compare these to two coverage-based, greedy SMBP baselines. The results reveal that the best SBSMBP technique is competitive with the best greedy baselines but does not improve on its effectiveness. Surprisingly, the Greedy technique relying only on the historical performance change size is sometimes more but at least equally effective as the best coverage-based, greedy baselines and SBSMBP techniques while being significantly more efficient.